# A New Causal Rule Learning Approach to Interpretable Estimation of Heterogeneous Treatment Effect

### v2

@@ 4.2 Real-world data application
In this part, we apply the CRL workflow to atrial septal defect (ASD) data to demonstrate its usage and performance (especially interpretability) in real-world observational data. ASD is a congenital heart defect characterized by a hole in the wall (atrial septum) that separates the upper chambers (atria) of the heart. As a common type of birth defect, ASD has a prevalence of 2.5 out of 1000 in live births and 25% to 30% present in adulthood [8, 43].
In this study, we focus on two common ASD treatments: percutaneous interventional closure (PIC) and minimally invasive surgical closure (MISC). PIC involves using catheters to reach the heart through blood vessels and repair the defect with an occluder, while MISC involves making a small incision in the chest, allowing access to the heart for patching. We investigate the HTE of PIC versus MISC (negative treatment) on hospital-free days within one year after discharge (HFD1Y). Hospital-free days are the days patients spend outside of the hospital. Instead of survival or alive days that overemphasize the survival goal, it is more pragmatic and patient-centered, reflecting more information on patient’s quality of life [6].
@@ 4.2.1 Overview of the ASD data
The ASD data used in this study is extracted from congenital heart disease data collected from thirteen clinical centers in China with the initiative to analyze and compare clinical pathways for first-surgery patients (patients undergoing surgery for the first time). The data includes 6780 patient records with 250 variables,spanning from hospital admission and preoperative examination to diagnoses, surgery, discharge, and follow-up visits. To implement our analysis, we extract the records of patients who were diagnosed with only ASD and treated with either PIC or MISC. There are only 1.16% missing values in these records, so we simply impute the data with K-Nearest Neighbor (number of neighbors set to 10). The final data includes 2850 observations, with 16 pre-treatment variables and five surgery-related variables, some of which are manually integrated with multiple original variables to make the variable set more concise. Pre-treatment variables include the information generated on admission, through the inquiry and diagnosis, and from pre-surgery medical check results. For instance, on admission, we have: i) two cumulative scores on disease history chdhis and othdhis. If a patient’s chdhis$=3$, it means that the patient has three previous cardiovascular conditions, ii) basic measurements with sysbp and diabp representing systolic and diastolic blood pressure respectively, and BMI the Body Mass Index, and iii) demographic information like age and sex.
We give details of the ASD data in Appendix D where Figure 1 shows the variable details and Table LABEL:tab:des.stat.realdata1 and Table 2 show the descriptive statistics for the pre-treatment variables and the surgery-related variables (and outcome HFD1Y), respectively.
Before we apply CRL to this observational data, we first need to pre-process the data using the propensity score matching method [49], which has proven to be very useful in estimating treatment effects using observational data [2]. This process allows us to match the data samples from both treatment groups to achieve pseudo randomization of treatment assignment and to overcome the unsatisfactory overlap between two groups, thus ensuring the satisfaction of the unconfoundedness assumption and overlap assumption. Additionally, this process models propensity scores that can be used in the D-learning step. Details of the matching process can be found in Appendix E, after which 26 observations are eliminated from the data and variable atsize0 is replaced by its square root, $sq.atsize0$ for better covariate balance. We then use the matched data from this process to fit all models for performance evaluation.
@@ 4.2.2 Performance metrics for real-world observational data
When estimating the treatment effect using real-world data, the ground truth is never known. Thus, we can no longer use the performance metrics defined for simulated data. Instead, we have to do so in some indirect ways based on what we know, like the following shown:
@@
ii) Empirical expected outcome (EEO). This metric is defined in [34] as:
@@
iii) Treatment efficient frontier, proposed in [47], is a graphic comparison of subgroups obtained from different methods, given that the identified subgroups differ in group size (proportion of samples) and CATE. Though not always the case, a small group usually has a higher CATE than a bigger one, so it is hard to tell which subgroup is better. To balance the trade-off between subgroup size and subgroup CATE, the authors define the concept of treatment efficient frontier, which is formed with all dominant groups (subgroup that has no other subgroup surpasses it in both size and CATE) a certain method identifies. This frontier is then mapped to the two-dimensional coordinate system with the $x$-axis being the size and $y$-axis the CATE estimates of the dominant groups. With this graph, we can see how quickly the estimated treatment effect magnitude decays as the subgroup covers more samples and easily spot the method that has the highest frontier since they find subgroups with larger sizes and CATE. See Figure 6 for example.
Figure 6: Treatment efficient frontiers of all methods fit for 100 repetitions
Table 1: Mean squared error (MSE) and empirical expected outcome (EEO) averaged on 100 repetitions of ASD data applying CRL and other baseline models. SD is the standard deviation of MSE and EEO.
@@
Table 2: Eight CRL subgroups that most impact the estimates of treatment effect, displayed in descending order of the absolute value of the subgroup CATE multiplied by the subgroup weight.
@@
$\times$×Weight
BMI$>$14.4 and electb=1 and oroom=1 and othchd$>$0 2.92 119.54 349.44
$\leq$9 and age$>$5 7.62 -42.42 -323.34
@@
$\leq$2 and sysbp$>$100 7.79 -31.14 -242.45
age$>$5 and BMI$>$17.2 and electb=1 6.09 -38.38 -233.79
age$>$15 and BMI$>$13.7 6.85 -33.52 -229.63
diabp$>$59 and electb=1 and oroom=0 6.80 29.20 198.66
Considering the above performance metrics, especially the ITE-based prediction accuracy, we implement all methods: We first split the ASD data into three parts: 40% (estimation set) for ITE estimation with all methods; 40% (prediction set) for outcome prediction using the least squares model (or any other proper prediction models) fed with all the same variables except the different ITE estimated by different methods in the previous step. The remaining 20% data (test set) is used to report the first two performance metrics, i.e., $MSE(g(\hat{\tau}(X))$ and EEO.
We repeat the above procedure 100 times with randomly re-sampled ASD data and tune the parameters using the same approach as in the simulation part. See Appendix C for details.
To evaluate treatment efficient frontier, we use the entire data to fit all models for each of the re-sampled data, since CRL generally needs more observations to discover potential subgroups. For each method to be compared, we pool the subgroups identified from all 100 repetitions together and find its dominant subgroups, which form the ultimate frontier.
4.2.3 Model results and interpretation
We list the MSE of the predicted HFD1Y and EEO of all methods averaged over 100 repetitions in the first and second rows in Table 1. It is obvious that all methods show similar accuracies of outcome prediction (MSE) and recommendation performance (EEO) on the ASD data. Figure 6 shows the treatment efficient frontier of all methods. Obviously, compared with other methods, CRL identifies subgroups with a variety of sizes ranging from 0.1 to 0.85 and effect sizes ranging from 4 to 25. In contrast, methods like OWE and CT, only identify big subgroups with relatively small subgroup CATE. Note that we do not show the frontier of CF in this figure as it overlaps with the frontier of CRL whose dominant groups are part of CF groups.
The above results demonstrate that CRL achieves at least comparable results against baseline methods in ITE estimation and subgroup identification on the ASD data. Moreover, it provides us with an enhanced understanding of ITE of ASD treatments on HFD1Y through the contributing informative rules we learned from the whole workflow.
For a concise presentation of what we learned from the ASD data with CRL, we list the eight rules that impact the estimated treatment effect the most in a randomly chosen rule set out of the 100 rule sets, with their corresponding CATE estimates and learned weights (See Table 2). The rules are displayed in decreasing order of the absolute value of the signed product of subgroup CATE multiplied by subgroup weight. For example, the first rule means that when a patient with abnormal ECG result, BMI$>$14.4, othchd$>$0 and is operated at a one-stop operating room, the patient’s ITE of PIC against MISC will increase by 349.44 days. However, this estimate will then decrease by 323.34 if the patient’s age is between five and nine. The final ITE estimate of a patient is the summation of all the products corresponding to the subgroups a patient belongs to, showing how a patient’s ITE is affected by the combination of potential rules.
These eight rules identify several important factors,such as age, BMI, electh, oroom, and NYHA, with meaningful cut-off value (e.g., 14.44 for BMI), influencing the treatment effect of PIC versus MISC on HFD1Y. Age and BMI are the two most involved factors in the rules, indicating they play indispensable roles in the way HFD1Y is affected by PIC and MISC. This is consistent with [23] and [33] stating that age and weight-related factors like BMI are particularly important in surgery decisions and the prognosis of ASD, such as length of hospital stay. Compared to MISC, PIC, in general, has the advantage of shorter hospital stay and is more suitable for non-underweight children or older children [33], which is reflected with the HFD1Y increase in the first (BMI$>$14.4) and second rule (age$>$8). For children aged between five and nine (second rule), the decrease of HFD1Y or the increase in hospital stay, may be due to a lower percentage of PIC-treated patients within the group compared to that of the remaining population (81% versus 85%). This reveals the preference for MISC against PIC in the young-aged group, which has been a fact in past ASD treatment practices [27].
In addition, these rules identify multiple cut-off values for the same factors, such as age with thresholds at 8 (third rule) and 15 (seventh rule), together with NYHA and sq.atsize0, revealing potential complex interactions between these factors. Although there is little evidence of these interactions, we believe they enlighten people in the community on new research directions, such as designing rigorous experiments to further validate these underlying relationships.
@@ 4.2.3 Model results and interpretation
othchd$>$0 and sq.atsize0$>$0 and $\leq$0 9.91 0.33
atrial$>$2 and $\leq$0 9.89 0.48
$\leq$2 and $\leq$0 9.79 0.56
[…]
atrial=2 7.68 0.86
[…]
We also list the details of the dominant subgroups of CRE (See Table 3) and PRIM (eight top rules by subgroup CATE, see Table 4) shown in Figure 6 for a better comparison on interpretability. Apparently, both CRE and PRIM identify relatively homogeneous rules that entail limited information. We find that for the same variable (except murmur), its cutoff value is almost the same in all the rules involved, such as 2 for atrial, NYHA and chdhis across all rules. Many rules share similar and even identical variables, such as atrial$>$2 and $\leq$0, $\leq$2, and $\leq$0 (the second and third rules) in Table 3 and the first two rules in Table 4, which further exacerbates this rule of homogenization. As a result, the subgroups identified are highly overlapping, hence the estimates of their treatment effects are also quite close. In contrast, CRL can explore a more diverse combination of variables for potential relationships in the ASD data. When different variables interact with each other, CRL identifies different cut-off values for each variable, which is more in line with the actual situation of complex disease, therefore providing much more interpretation and insights to healthcare practitioners.
Subgroup CATE Size
[…]
and $\leq$2 and $\leq$2 and $\leq$0 10.46 0.48
and $\leq$5 and $\leq$1 and $\geq$3
[…]
$\leq$1 and $\leq$1 and $\leq$2 9.96 0.52
and $\leq$0 and $\leq$0 and $\leq$0 and $\geq$2
[…]
$\leq$0 and $\leq$1 and $\geq$2
[…]
$\leq$1 and $\leq$2 and $\leq$5 8.68 0.69
and $\leq$0 and $\leq$2 and $\leq$0
@@ 5 Conclusions and future work
This study introduces a rule-based framework, CRL, aimed at enhancing our understanding of HTE estimation for complex disease treatment in clinical scenarios. CRL leverages causal forest to discover a pool of potential rules for ITE estimation (rule discovery) and then utilizes the D-learning method to filter out non-informative rules, imposing sparsity on the rules (rule selection). By deconstructing the estimated ITE through a weighted combination of the informative rules selected by D-learning, CRL explores HTE with a novel perspective that connects group-level and individual-level treatment effects. The understandable rules, their corresponding CATE estimates, weights in combination, and the comprehensive procedure for evaluating each causal rule from multiple perspectives (rule analysis) together give insights into the treatment of a complicated disease. The remaining rules that survive after the rule analysis step may reveal potential meaningful interactions between involved variables, paving the way for further investigation and validation.
This research has several methodological limitations that could be addressed in future studies. First of all, CRL requires the outcome of interest be continuous and prefer higher values. Treatment in our study is also limited to the binary case. These limits hinder the wide application of the method in healthcare practice. Secondly, our assumption of ITE is based on a linear combination of potential rules for simplicity, while the non-linear case is not discussed. Last but not least, in the evaluation of CRL from treatment recommendation, we simply recommend treatments based on the sign of the estimated ITE and do not consider the cost-sensitive scenarios where treatment resource are constrained. Future work may introduce more economic factors and explore how to better estimate treatment effects and recommend treatments in complex scenarios.
@@
The data that support the findings of this study contain sensitive patient information and, therefore, cannot be made publicly available. However, these data are available from the corresponding author upon reasonable request, subject to appropriate confidentiality agreements.

### v3

@@ 4.4 Real-world data application
In this part, we apply the CRL workflow to atrial septal defect (ASD) data to demonstrate its usage and performance (especially interpretability) in real-world observational data. ASD is a congenital heart defect characterized by a hole in the wall (atrial septum) that separates the upper chambers (atria) of the heart. As a common type of birth defect, ASD has a prevalence of 2.5 out of 1000 in live births and 25% to 30% present in adulthood [brida2021atrial, vanderlinde2011birth].
In this study, we focus on two common ASD treatments: percutaneous interventional closure (PIC) and minimally invasive surgical closure (MISC). PIC involves using catheters to reach the heart through blood vessels and repair the defect with an occluder, while MISC involves making a small incision in the chest, allowing access to the heart for patching. We investigate the HTE of PIC versus MISC (negative treatment) on hospital-free days within one year after discharge (HFD1Y). Hospital-free days are the days patients spend outside the hospital. Instead of survival or days alive that overemphasize the survival goal, this metric is more pragmatic and patient-centered, reflecting more information on the patient’s quality of life [auriemma2021hospital].
@@ 4.4.1 Overview of the ASD data
The ASD data used in this study were extracted from congenital heart disease data collected from thirteen clinical centers in China, with the initiative to analyze and compare clinical pathways for first-surgery patients (patients undergoing surgery for the first time). The data includes 6780 patient observations with 250 variables, spanning from hospital admission and preoperative examination to diagnoses, surgery, discharge, and follow-up visits. To implement our analysis, we extract the observations of patients who were diagnosed with only ASD and treated with either PIC or MISC. There are only 1.16% missing values in these observations, so we simply impute the data with K-Nearest Neighbor (number of neighbors set to 10). The final data includes 2850 observations, with 16 pre-treatment variables and five surgery-related variables, some of which are manually integrated with multiple original variables to make the variable set more concise. Pre-treatment variables include the information generated on admission, through the inquiry and diagnosis, and from pre-surgery medical check results. For instance, on admission, we have: i) two cumulative scores on disease history chdhis and othdhis. If a patient’s chdhis=3=3, it means that the patient has three previous cardiovascular conditions, ii) basic measurements with sysbp and diabp representing systolic and diastolic blood pressure, respectively, and BMI the Body Mass Index, and iii) demographic information like age and sex.
We give details of the ASD data in Appendix F, where Figure 1 shows the variable details and Table LABEL:tab:des.stat.realdata1 and Table 2 show the descriptive statistics for the pre-treatment variables and the surgery-related variables (and outcome HFD1Y), respectively.
Before we apply CRL to this observational data, we first need to pre-process the data using the propensity score matching method [zhao2021propensity], which has proven to be very useful in estimating treatment effects using observational data [Psmcomparison]. This process allows us to match the data samples from both treatment groups to achieve pseudo-randomization of treatment assignment and to overcome the unsatisfactory overlap between the two groups, thus ensuring the satisfaction of the unconfoundedness assumption and overlap assumption. Additionally, this process models propensity scores that can be used in the D-learning step. Details of the matching process can be found in Appendix G, after which 26 observations are eliminated from the data and variable atsize0 is replaced by its square root, $sq.atsize0$, for better covariate balance. We then use the matched data from this process to fit all models for performance evaluation.
@@ 4.4.2 Performance metrics for real-world observational data
When estimating the treatment effect using real-world data, the ground truth is never known. Thus, we can no longer use the performance metrics defined for simulated data. Instead, we have to do so in some indirect ways based on what we know, like the following:
@@
ii) Empirical expected outcome (EEO). This metric is defined in [qi2018d] as:
@@
iii) Mean-squared prediction error (MSPE) of transformed outcome. Proposed in [hitsch2024heterogeneous] as $\operatorname{MSPE}=\mathbb{E}[(Y_{i}^{*}-\hat{\tau}(X_{i}))^{2}]$. The transformed outcome can be calculated from the observed outcome $Y_{i}$ and propensity score $\pi(1,X)$:
$Y_{i}^{*}=\frac{A_{i}+1}{2}\cdot\frac{Y_{i}}{\pi(1,X_{i})}-\frac{1-A_{i}}{2}\cdot\frac{Y_{i}}{1-\pi(1,X_{i})}.$
$Y_{i}^{*}$ is a well-established unbiased estimate of $\tau(X_{i})$ under the unconfoundedness assumption, i.e., $\mathbb{E}[Y_{i}^{*}|X_{i}]=\tau(X_{i})$. Consequently, the MSPE serves as a proxy for the unobservable mean-squared error $\mathbb{E}[(\tau(X_{i})-\hat{\tau}(X_{i}))^{2}]$. Therefore, an estimator with a lower MSPE is considered to provide a more accurate estimate of the true treatment effect.
iv) Treatment efficient frontier, proposed in [wang2021causal], is a graphic comparison of subgroups obtained from different methods, given that the identified subgroups differ in group size (proportion of samples) and CATE. Though not always the case, a small group usually has a higher CATE than a bigger one, so it is hard to tell which subgroup is better. To balance the trade-off between subgroup size and subgroup CATE, the authors define the concept of treatment efficient frontier, which is formed with all dominant groups (subgroup that has no other subgroup surpassing it in both size and CATE) a certain method identifies. This frontier is then mapped to the two-dimensional coordinate system with the $x$-axis being the size and $y$-axis the CATE estimates of the dominant groups. With this graph, we can see how quickly the estimated treatment effect magnitude decays as the subgroup covers more samples and easily spot the method that has the highest frontier since it finds subgroups with larger sizes and CATE. See Figure 8 for example.
Figure 8: Treatment efficient frontiers of all methods fit for 100 repetitions.
Table 2: Mean squared error (MSE) of outcome prediction and empirical expected outcome (EEO) averaged over 100 repetitions of ASD data applying CRL and other baseline models. SD is the standard deviation of MSE and EEO.
@@
Considering the above performance metrics, especially the ITE-based prediction accuracy, we implement all methods: We first split the ASD data into three parts: 40% (estimation set) for ITE estimation with all methods; 40% (prediction set) for outcome prediction using the least squares model (or any other proper prediction models) fed with all the same variables except the different ITE estimated by different methods in the previous step. The remaining 20% data (test set) is used to report the first two performance metrics, i.e., $MSE(g(\hat{\tau}(X))$ and EEO. We repeat the above procedure 100 times with randomly re-sampled ASD data.
To compute the MSPE, we employ a bootstrapped sampling approach: we split our ASD data into a training set (70% samples) for fitting the estimation model and a test set (30% samples) for reporting MSPE. We then draw 1,000 bootstrapped samples with replacement from each dataset, ensuring no data sample is contained in both the training and test sets. The MSPE values reported represent the average across these 1000 bootstrap samples. In both the application of CRL to compute the above three metrics, we tune the relevant parameters using the same approach as in the simulation part. See Appendix D for details.
To evaluate the treatment efficient frontier, we use the entire data to fit all models for each of the resampled data, since CRL generally needs more observations to discover potential subgroups. For each method to be compared, we pool the subgroups identified from all 100 repetitions together and find its dominant subgroups, which form the ultimate frontier.
4.4.3 Model results and interpretation
Table 3: Mean-squared prediction error (MSPE) of transformed outcome with standard error (SE) and standard deviation (SD) across the 1000 bootstrap test samples of ASD data applying CRL and other baseline models.
Mean MSPE SE SD
CRL 81775.58 270.61 8557.44
CF 1184278 12962.8 409919.7
CT 1183005 12975.51 410321.8
OWE 1166325 13303.69 420699.7
PRIM 1187621 13147.51 415760.9
CRE 1184111 12961.91 409891.7
We list the MSE of the predicted HFD1Y and EEO of all methods averaged over 100 repetitions in the first and second rows in Table 2. It is obvious that all methods show similar accuracies of outcome prediction (MSE) and recommendation performance (EEO) on the ASD data. The summary statistics of MSPE are shown in 3 where CRL demonstrates dramatically lower values compared to all baseline methods. Figure 8 shows the treatment efficient frontier of all methods. Obviously, compared with other methods (except CF), CRL identifies subgroups with a variety of sizes ranging from 0.1 to 0.85 and effect sizes ranging from 4 to 25. In contrast, methods like OWE and CT only identify big subgroups with relatively small subgroup CATE. The overlap between the CF and CRL frontiers is primarily driven by the CRL methodology itself: since CRL selects its rules from candidates generated by CF, its frontier inherently constitutes a refined, elite subset of the original CF rules, even after substantial filtering. Crucially, this demonstrates CRL’s ability to achieve predictive performance comparable to CF while yielding a more parsimonious, interpretable, and clinically actionable set of subgroups. The above results all demonstrate the effectiveness and superiority of CRL compared to other baseline methods on estimating HTE accurately. Moreover, it provides us with an enhanced understanding of ITE of ASD treatments on HFD1Y through the informative rules we learned from the whole workflow.
For a concise presentation of what we learned from the ASD data with CRL, we list the eight rules that impact the estimated treatment effect the most in a randomly chosen rule set out of the 100 rule sets, with their corresponding CATE estimates and learned weights (See Table 4). The rules are displayed in decreasing order of the absolute value of the signed product of subgroup CATE multiplied by subgroup weight. For example, the first rule means that when a patient with an abnormal ECG result, BMI>>14.4, othchd>>0, and is operated at a one-stop operating room, the patient’s ITE of PIC against MISC will increase by 349.44 days. However, this estimate will then decrease by 323.34 if the patient’s age is between five and nine. The final ITE estimate of a patient is the summation of all the products corresponding to the subgroups a patient belongs to, showing how a patient’s ITE is affected by the combination of potential rules.
These eight rules identify several important factors, such as age, BMI, electb, oroom, and NYHA, with meaningful cut-off values (e.g., 14.44 for BMI), influencing the treatment effect of PIC versus MISC on HFD1Y. Age and BMI are the two most involved factors in the rules, indicating they play indispensable roles in the way HFD1Y is affected by PIC and MISC. This is consistent with [hughes2002prospective] and [qi2020open] stating that age and weight-related factors like BMI are particularly important in surgery decisions and the prognosis of ASD, such as length of hospital stay. Compared to MISC, PIC, in general, has the advantage of shorter hospital stay and is more suitable for non-underweight children or older children [qi2020open], which is reflected with the HFD1Y increase in the first (BMI>>14.4) and second rule (age>>8). For children aged between five and nine (second rule), the decrease in HFD1Y or the increase in hospital stay may be due to a lower percentage of PIC-treated patients within the group compared to that of the remaining population (81% versus 85%). This reveals the preference for MISC over PIC in the young-age group, which has been a fact in past ASD treatment practice [karamlou2008rush].
@@ 4.4.3 Model results and interpretation
Table 4: Eight CRL subgroups that have the greatest impact on the estimates of treatment effect, displayed in descending order of the absolute value of the subgroup CATE multiplied by the subgroup weight.
@@
[…]
BMI>>14.4 and electb=1 and oroom=1 and othchd>>0 2.92 119.54 349.44
@@
[…]
$mur\leq$2 and sysbp>>100 7.79 -31.14 -242.45
age>>5 and BMI>>17.2 and electb=1 6.09 -38.38 -233.79
age>>15 and BMI>>13.7 6.85 -33.52 -229.63
diabp>>59 and electb=1 and oroom=0 6.80 29.20 198.66
[…]
@@ 5 Conclusions and future work
This study introduces a rule-based framework, CRL, aimed at enhancing our understanding of HTE estimation for complex disease treatment in clinical scenarios. CRL leverages the causal forest to discover a pool of potential rules for ITE estimation (rule discovery) and then utilizes the D-learning method to filter out non-informative rules, imposing sparsity on the rules (rule selection). By deconstructing the estimated ITE through a weighted combination of the informative rules selected by D-learning, CRL explores HTE with a novel perspective that connects group-level and individual-level treatment effects. The understandable rules, their corresponding CATE estimates, weights in combination, and the comprehensive procedure for evaluating each causal rule from multiple perspectives (rule analysis) together give insights into the treatment of a complicated disease. The remaining rules that survive after the rule analysis step may reveal potential meaningful interactions between involved variables, paving the way for further investigation and validation.
Subgroup CATE Size
[…]
This research has several methodological limitations that could be addressed in future studies. First of all, CRL requires the outcome of interest to be continuous and prefers higher values. Treatment in our study is also limited to the binary case. These limits hinder the wide application of the method in healthcare practice. Secondly, our assumption of ITE is based on a linear combination of potential rules for the simplicity of demonstration, while the non-linear case is not discussed and evaluated. Thirdly, in the evaluation of CRL from treatment recommendation, we simply recommend treatments based on the sign of the estimated ITE and do not consider the cost-sensitive scenarios where treatment resources are constrained. Future work may introduce more economic factors and explore how to better estimate treatment effects and recommend treatments in complex scenarios. Lastly, while the current staged architecture of CRL achieves superior interpretability and clinical practicality, it may result in lower efficiency compared to a fully integrated framework. We will explore the possibility of a more unified approach that maintains interpretability while improving operational efficiency in future work.
@@
The real-world data that support the findings of this study contain sensitive patient information and, therefore, cannot be made publicly available. However, these data are available from the corresponding author upon reasonable request, subject to appropriate confidentiality agreements.